OpenStudy (tray):

June bought ​ 3 1/2 ​ gallons of white paint at $17.50 per gallon and ​ 1 3/4 ​gallons of blue paint at $19.00 per gallon. What was her total cost?

OpenStudy (wanell):

312* 17.5= 5460.00 134*19=2546 2546+5460= $8,006.00

OpenStudy (tray):

8,006.00 ???

OpenStudy (will.h):

Not correct

OpenStudy (tray):

fractions?

OpenStudy (tray):